log(7-2x)=log(x+5)
Answer by fractalier(2101)

Since log(7 - 2x) = log(x + 5), we can exponentiate both sides, thus putting the arguments equal...thus we have
7 - 2x = x + 5 and
3x = 2 and
x = 2/3
Check to be sure this is a legal solution...it is, so
x = 2/3